1. Using standard heats of formation, calculate the standard enthalpy change for the following reaction.
2SO2(g) + O2(g)2SO3(g)
ANSWER: ___kJ
2. Using standard heats of formation, calculate the standard enthalpy change for the following reaction.
2H2O(l)2H2(g) + O2(g)
ANSWER: ___kJ
1
Expert's answer
2019-10-31T07:29:37-0400
  1. The standard enthalpy change of formation for compounds are

"\Delta"Hf0 SO2 = -296.84 kJ/mol

"\Delta" Hf0 SO3 = -395.7 kJ/mol

"\Delta" Hf0 O2 = 0 kJ/mol

So, the enthalpy change of reaction will be

"\Delta" Hrxn = [2moles*(-395.7kJ/mol)] - [2moles*(-296.84kJ/mol) + 1mole*0kJ/mol] = -791.4 kJ + 593.68 kJ = -197.7 kJ

2) The products are elements, so the enthalpy change forse them is 0.

For water is -285,8 for one mole, but there are 2:

"\Delta"Hrxn = 0-(-285,8x2) = +572 kJ